# Quasiregular maps of Sierpiński carpet Julia sets

### v1

@@ 1. Introduction
Acknowledgements. The authors thank the Institute for Mathematical Sciences at Stony Brook University for its hospitality. The second author also thanks the University of Science and Technology of China for the support of the travel that resulted in this work.
@@ 8. Quasiregular maps between other topological types of Julia sets
There exists a postcritically finite rational map $f\colon\mathbb{C}\to\mathbb{C}$ with gasket Julia set $\mathcal{J}(f)$, and a global quasiregular map $\xi\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ with $\xi^{-1}(\mathcal{J}(f))=\mathcal{J}(f)$, such that $\xi|_{\mathcal{J}(f)}$ does not have a rational extension to $\widehat{\mathbb{C}}$ satisfying Equation (2).
@@
Suppose that $\xi|_{\mathcal{J}(f)}$ can be extended to a global rational map $\widehat{\xi}$ satisfying Equation (2), i.e.,
$f^{k}\circ(\widehat{\xi})^{l}=f^{2k},$
for some $k,l\in\mathbb{N}$. However, in $\Delta ABC$, the left side of this equation is $f^{k+2l}$, and outside of this triangle it is $f^{k}$. But the right side is always $f^{2k}$, a rational map, a contradiction.

### v2

@@ 1. Introduction
Acknowledgements. The authors thank the Institute for Mathematical Sciences at Stony Brook University for its hospitality. The second author thanks the University of Science and Technology of China for the support of the travel that resulted in this work. The authors also thank Yusheng Luo for useful comments and references.
@@ 8. Quasiregular maps between other topological types of Julia sets
There exists a postcritically finite rational map $f\colon\mathbb{C}\to\mathbb{C}$ with gasket Julia set $\mathcal{J}(f)$, and a global quasiregular map $\xi\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ with $\xi^{-1}(\mathcal{J}(f))=\mathcal{J}(f)$, such that $\xi|_{\mathcal{J}(f)}$ does not have a rational extension to $\widehat{\mathbb{C}}$.
@@
Finally, $\xi|_{\mathcal{J}(f)}$ cannot be extended to a global rational map. Indeed, if it extended to a rational map of $\widehat{\mathbb{C}}$, then, by uniqueness for analytic maps, it would have to equal to $f^{2}$ in $\Delta ABC$ and to the identity in $\widehat{\mathbb{C}}\backslash\left((\Delta ABC)\cup\sqcup_{i=1}^{3}D_{i}\right)$. This is a contradiction since $f^{2}$ is not equal to the identity.